OpenStudy (anonymous):

find cot theta if (-6,-8) is a point on the terminal side of theta

OpenStudy (anonymous):

the reciprocal of tangent. or if the terminal side contains (a,b) then \[\cot(\theta)=\frac{a}{b}\]

OpenStudy (amistre64):

the cotangent function is simply defined as x/y , or a/b as the case may be :)

OpenStudy (anonymous):

meaning the answer in this case is -6/-8?

OpenStudy (amistre64):

yes, but reduce it as wanted

OpenStudy (anonymous):

yes satellite is absolutely right cot(theta) = a/b

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

i think that , any other problem